A Business Management monthly test was administered to a class consisting of 40 students. The mean score was 70 marks with a standard deviation of 6. Calculate the probability that a student obtained this score:

More than 82 marks
Between 64 and 72 marks
Less than 50 marks
1
Expert's answer
2019-08-19T13:31:55-0400

Define mean score m = 70, standard deviation "\sigma=6", score obtained by a student x.

Probability "P(a<x<b)=P(\frac{a-m}{\sigma}<z<\frac{b-m}{\sigma})", where z has standard normal distribution.

"P(82<x<\infty)=P(\frac{82-70}{6}<z<\infty)=\\
P(2<z<\infty)=1-P(z\leq2)=1-0.9772=0.0228"

"P(64<x<72)=P(\frac{64-70}{6}<z<\frac{72-70}{6})=\\
P(-1<z<1/3)=P(z< 1/3)-P(z\leq-1)=\\
0.6306-0.1587=0.4719"

"P(-\infty<x<50)=P(-\infty<z<\frac{50-70}{6})=\\
P(-\infty<z<-20/6)=0.0004"

Answer: P(x > 82) = 0.0228, P(64 < x <72) = 0.4719, P(x < 50) = 0.0004.